OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations auditioned today?

OpenStudy (anonymous):

430 720 1150 1300

OpenStudy (campbell_st):

a = 1 step equations b = 2 step equations a+ b = 1150 (1) 3a -2b = 1300 (2) equation (1) x 2 + equation (2) will eliminate b 5a = 3600 a = 720 then b = 430 there were 720 one step equations and 430 two step equations